As we know that maximum recombinant frequency possible is 50% & in case of linked genes it is less than 50. But in Mendel's experiment in F2 (dihybrid) the recombinant frequency was 37.5% even though the genes were unlinked( on different chromosomes). Why it comes 37.5% and not 50%?

1
Expert's answer
2022-05-16T17:20:01-0400

In dihybrid cross, the recombination frequency is 37.5% for the following reason.


Each of the offspring of the first generation produces 4 different gametes, 2 of which are recombinant, that is, not similar to the parent ones. Accordingly, in a dihybrid cross, the frequency of occurrence of recombinant gametes is 50%, as in the case of a monohybrid cross, in which 1 recombinant gamete occurs out of all 2 gametes.


The probability that in a dihybrid cross one recombinant gamete of a certain type will merge with the same type is:

0.25*0.25 = 0.0625.

Since there are 2 recombinant gametes in total, then:

2*0.0625 = 0.125.


The combination of two different recombinant gametes is not taken into account, since in this case in the second generation a genotype will be released that expresses the parental phenotype.


Not only recombinant gametes, but also one of the parental gametes, which contains only recessive alleles, take part in the compilation of recombinant genotypes. A parent gamete that contains only dominant alleles, when matched with any other gamete, will always give a genotype that expresses the parental phenotype, which means that we do not take such a gamete into account.


A parent gamete that contains only recessive alleles, when copulated with recombinant gametes, also produce a recombinant genotype. But we do not consider the copulation of such a parental gamete with itself, since in this case the parental genotype will be released in the second generation.


The probability that each of two recombinant gametes in one offspring of the first generation will fuse with a parent gamete that contains only recessive alleles is:

2*0.25*0.25 = 0.125.


But the same happens in another offspring of the first generation, which means that the probability of complete copulation of recombinant gametes with parental gametes that contain only recessive alleles is:

2*0.125 = 0.25.


If you add one value in bold above with another of the same value, then:

0.125 + 0.25 = 0.375 = 37.5%.